can someone please tell me if these are correct, thank you
T=v^2p^2/c if v=3, p=2 and c=4
MY ANSWER T=50
T=xh^2 if x=2 and h=5
MY ANSWER T=18

Answer by checkley71(8403)

T=3*3*2*2/4 OR T=36/4 OR T=9 [V~2=2*2 & P~2=3*3]
T=2*5*5 OR T=50 [H~2=5*5]